The very small, frequent earthquakes and continuous broad-band tremor recorded at the Gages seismic station since 28 November have continued. The amplitude of the tremor is still increasing slightly, and it is now being recorded at all the stations in the network, though at very low amplitudes at most stations. Five (5) earthquakes have been located for this period, at depths of less than 2 km. Four of these earthquakes occurred beneath the volcano, and one was located near Harris'.
All of the seven EDM lines have been remeasured today. Once again, no significant changes were recorded in any of the line lengths. GPS measurements were made overnight. The data are still being processed so results are not yet available.
Visual observations of the crater area were made from the helicopter this morning. There was vigorous steam production from several areas, and continued slow growth of the lava dome within the July 18 vent was noted. It was confirmed that a second lava dome is actively growing from an area to the north of the September spine; the rate of growth of this dome is also thought to be low. Small ash clouds were observed from the crater area at 11:30 and 14:30.
Some changes to the staff at the observatory have occurred during this period. Dr Ellen Howell has left the island and has been replaced by Don Hooper, who will continue the GPS monitoring work. Richie Robertson returned to the island from SRU headquarters in Trinidad yesterday evening to rejoin the monitoring team after a short absence.